1. From the following information, calculate the volatility of the portfolio and comment

on the relationship among the stocks return.

Day Stock A Stock B Stock C 

1 0.4      2.2      0.6

2 1.1      1.3      0.5 

3 0.9      1.2      1.4

4 1.7      1.9     1.6


1
Expert's answer
2021-10-31T18:27:27-0400

portfolio volatility = "\sqrt{Var(X+Y+Z)}"

"Var(X+Y+Z)=Var(X)+Var(Y)+Var(Z)"

X - Stock A, Y - Stock B, Z - Stock C

"Var(X)=0.217,Var(Y)=0.1725,Var(Z)=0.232"


portfolio volatility = "\sqrt{0.217+0.1725+0.232}=0.788"


relationship among the stocks return:

correlation coefficients:

"r_{XY}=-0.187" - a non significant very small negative relationship

"r_{XZ}=0.633" - a non significant large positive relationship

"r_{YZ}=-0.094" - a non significant very small negative relationship